OpenStudy (anonymous):

Evaluate the radical expression and express the result in the form a + bi. Square root of −81

OpenStudy (zarkon):

9i=0+9i

OpenStudy (anonymous):

how did you do that so fast

OpenStudy (zarkon):

\[\sqrt{81}=9\] \[\sqrt{-1}=i\]